### v4

@@ 2 Fractionalized Fermi liquids
An important step forward was Oshikawa’s proof of the Luttinger constraint using a ’t Hooft anomaly matching argument [123]. Oshikawa identified a mixed anomaly between translations and global U(1) symmetry of charge conservation. The value of the anomaly can be computed exactly from a knowledge of the density of electrons in the lattice Hamiltonian. Matching this anomaly with that of the low energy Fermi liquid theory, Oshikawa established the Luttinger constraint in a non-perturbative manner [40, 124, 4].
Soon after, the idea of a ‘fractionalized Fermi liquid’ (FL*) was introduced [36, 37], as a metallic state in which the Fermi surface did not obey the Luttinger constraint. In the simplest case, the volume enclosed by the Fermi surface in FL* was the same as free electrons with density $\rho-1$ (we are assuming a spin $S=1/2$ degeneracy). The central point was that it was possible to satisfy Oshikawa’s anomaly by combining the anomaly of a Fermi surface (which contributes an amount equivalent to a density $\rho-1$) with that of a fractionalized spin liquid of the type we will study in Section 3 (which contributes an amount equivalent to a density 1). It is trivially possible to shift $\rho$ by an even integer by adding and removing filled bands, and the novelty is the shift in FL* by an odd integer.
@@
These ideas apply rather generally to electronic lattice models, and are summarized schematically in Fig. 3. An explicit construction of FL* is simplest for the Kondo lattice model, and this will be described in Section 2.1. Our main interest here is the FL* phase in a single band Hubbard-like model, and we will turn to this in Section 2.2.
@@ 2.1 Kondo lattice
The $\Phi$ is the decoupling field for the $J_{K}$ interaction, while the $t_{1,{\bm{i}}{\bm{j}}}$ are the decoupling fields for the $J_{1,{\bm{i}}{\bm{j}}}$ interactions. Unlike Section 3.2, we do not assume any flux in the $t_{1,{\bm{i}}{\bm{j}}}$, and make the simplest possible choice in which the $t_{1,{\bm{i}}{\bm{j}}}$ have the same symmetry as the lattice. With these choices, our Kondo lattice theory only has an emergent U(1) gauge symmetry after adding an emergent gauge field $a_{{\bm{i}}{\bm{j}}}$ via the Peierls factor $t_{1,{\bm{i}}{\bm{j}}}\rightarrow t_{1,{\bm{i}}{\bm{j}}}e^{ia_{{\bm{i}}{\bm{j}}}}$. The gauge transformations are
@@
  • $\langle\Phi\rangle=0$, FL*. This is the novel phase, in which the $f_{1}$ and $c$ fermions are decoupled to leading order in $\mathcal{H}_{\rm KLmf}$. The $f_{1\alpha}$ form a spin liquid, in this case one with a spinon Fermi surface. But any other spin liquid with the same anomaly is allowed i.e. any other spin liquid with unit density of spinons in the ground state. The conduction electrons now form a ‘small’ Fermi surface of area $(\rho-1)/2=p/2$ (mod 1) per spin (where $\rho$ is the total density of electrons of both spins in the Kondo lattice), and this is not the Luttinger value. At higher order, the $f_{1}$ and $c$ fermions will couple with each other, but by extending the general arguments of Luttinger, we can conclude that such effects will not charge the Fermi surface area. Note that a solution with $\langle\Phi\rangle=0$ is not permitted in the single-spin Kondo model because $J_{K}$ flows to infinity under renormalization. In the lattice model, the flow of $J_{K}$ is cutoff at a scale determined by $J_{1{\bm{i}}{\bm{j}}}$ [125].
@@
  • FL* in the Anderson lattice. It also interesting to consider the formation of FL* in an Anderson lattice model, where the ${\bm{S}}_{1}$ spins are replaced by electronic orbitals with a large local repulsion $U$, and the total density of electrons in both bands remains at $1+p$. In general, the Anderson lattice model is not particle-hole symmetric, and so the density of electrons in the $f_{1}$ band will deviate from unity, and consequently the actual density of electrons in the $c$ band will deviate from $p$. Nevertheless, a crucial point is that the area of the Fermi surface per spin will remain pinned at $p/2$ (mod 1) because the anomaly of the spin liquid is pinned at unity.
@@ 2.2 Single band model
We now turn to a construction of the FL* phase in a single band model, such as the square lattice Hubbard model (possibly with additional short-range interactions) of interest for the cuprates. Now the situation is more complicated than in the Kondo lattice model, as there is no natural distinction between electrons that form a spin liquid, and electrons that form a Fermi surface.
Nevertheless, at the level of cartoon pictures, we can illustrate the structure of a FL* state in Fig. 6 [126]. This figure describes three distinct metallic phases in a Hubbard time model with electron density 1p1-p. Recall that in the absence of a broken symmetry, the Luttinger constraint on hole Fermi surfaces is that they have a fractional area per spin of $(1+p)/2$, relative to the area of the full square lattice Brillouin zone.
Figure 6: Cartoon pictures of different states of doped antiferromagnets; adapted from Ref. [126]. The areas are those that would be measured by a probe such as quantum oscillation. The AF metal has long-range antiferromagnetic order, and the reduced Brillouin zone is shown with the dashed line. The other phases do not break any symmetries. The open circles are holons, and these are assumed fermionic in the holon metal. The green dimers represent bound states of holons and spinons. Note that all 3 states would show a Hall co-efficient of density $p$ positively charged carries.
@@
  • Holon metal. This is a state with no broken symmetry, in which the electrons have paired up in singlet bonds which resonate with each other. The dopants are realized by spinless mobile vacancies of charge +e+e, known as holons. The density of holons is $p$, and if the holons are fermions, they will form Fermi surfaces corresponding to spinless free fermions of density $p$. If there are four distinct Fermi surfaces in the Brillouin zone (as is the case in many computations), then the fractional area of each pocket will be $p/4$. Although this area is the same as that for the AF metal, the reason is very different. Now there is no broken symmetry, and the fermionic quasiparticles are spinless holons. This Fermi surface area does not obey the Luttinger constraint, and this is permitted because of the presence of the spin liquid. A systematic, gauge theoretic treatment of the holon metal state is available in the literature [127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137]. We note in passing that with the generalized definition in Fig. 3, the holon metal is also a type of FL*.
  • FL*. Finally, we turn to the metallic state of interest. The holon metal state also has spinon excitations, which can be created out of the ground state. Now imagine a situation in which each holons gains energy by binding with a spinon, so that the system can pay the price for creating the spinons. Then the ground state will change into one in which the mobile charge carriers are holon-spinon bound states [138, 52, 139, 140, 141, 142, 143, 144, 145, 126, 146, 147, 148, 149, 150]. These bound states are always fermions with charge +e+e, spin $S=1/2$, just like a hole. Treating these holes as free fermions, we conclude that the total area of the Fermi surface should be $p/2$. If there are 4 distinct pockets (as there in the computation below), then each pocket will have the distinctive area of $p/8$. This Fermi surface area also does not obey the Luttinger constraint.
As we noted in Section 1.1, and will discuss further in Section 5, recent observations by Chan et al. [18] of the Yamaji effect in the cuprate HgBa2CuO4+δ show a fractional area of ‘approximately $1.3\%$’ at a doping $p=0.1$, close to the value $p/8=1.25\%$ predicted for FL* [36, 37, 52, 151]. This is strong evidence in favor a FL* description of the pseudogap phase of the cuprates.
@@ 2.2.1 Layer construction with ancilla qubits
We wish to have a mean-field theory which changes a large hole-like Fermi surface of area $(1+p)/2$ to small hole-like Fermi surfaces of total area $p/2$. The simplest way to achieve this in mean-field theory is to hybridize the large Fermi surface with another band at half-filling (as in the Kondo lattice)—this leads to a Fermi surface of the needed area $(1+p+1)/2$ (mod 1) =p/2=p/2. But we are not allowed to add a single band at half-filling because it is not ‘trivial’ i.e. it is not anomaly-free, and its excitations cannot be integrated out because the extra band can only acquire gap with a broken symmetry. On the other hand, we are allowed to add two bands at half filling each (or any even number of bands), because they can form a trivial insulator with a gap. After the first added band hybridizes with the physical electrons to yield the small Fermi surfaces, the second added band is left decoupled, and it must form a spin liquid to satisfy the Oshikawa anomaly. This is the essence of the ALM.
The method is illustrated in Fig. 7. We begin with a Hubbard model and augment it with ancillas which are decoupled from the Hubbard model.
Figure 7: Illustration of the mapping from a single-band Hubbard model with decoupled ancilla qubits, to a single band with free fermions coupled to a bilayer antiferromagnet. The Schrieffer-Wolff transformation is derived in Ref. [152], and the Hubbard-Stratonovich transformation is derived in Ref. [51]. The layer numbers of the layer construction with ancilla qubits are indicated in the bottom picture; these appear in Tables 2 and 3.
It is important that the new degrees of freedom have a trivial ground state with an energy gap, so that its extra excited states can be eliminated by a canonical transformation. For the ancillas, we choose a pair of qubits associated with every site of the square lattice, which we identify as $S=1/2$ spins ${\bm{S}}_{1{\bm{i}}}$ and ${\bm{S}}_{2{\bm{i}}}$. There is an antiferromagnetic exchange interaction $J_{\perp}$ between the two spins at each $\bm{i}$}, and so the ancillas form a bilayer square lattice antiferromagnet. We take $J_{\perp}$ to be large enough so that the ground state of the bilayer is smoothly connected to the ‘trivial’ rung singlet state. The decoupled ancilla Hamiltonian is therefore
$\mathcal{H}_{\rm ancilla-decoupled}$ $=$ $-\sum_{{\bm{i}},{\bm{j}}}t_{{\bm{i}}{\bm{j}}}c^{\dagger}_{{\bm{i}}\alpha}c_{{\bm{j}}\alpha}+U\sum_{\bm{i}}(c_{{\bm{i}}\uparrow}^{\dagger}c_{{\bm{i}}\uparrow})(c_{{\bm{i}}\downarrow}^{\dagger}c_{{\bm{i}}\downarrow})$ (2.6)
$~~+J_{\perp}\sum_{{\bm{i}}}{\bm{S}}_{1{\bm{i}}}\cdot{\bm{S}}_{2{\bm{i}}}+\ldots\,$,
where the ellipses represent exchange interactions within the ancilla layers.
The Ancilla Layer Model Hamiltonian, $\mathcal{H}_{\rm ALM}$, we shall work with couples the electronic sector to the ancillas. It is a simple augmentation of the Kondo lattice Hamiltonian $\mathcal{H}_{\rm KL}$ in Eq. (2.1) by the layer of ${\bm{S}}_{2{\bm{i}}}$ spins
$\mathcal{H}_{\rm ALM}=\mathcal{H}_{\rm KL}+J_{\perp}\sum_{{\bm{i}}}{\bm{S}}_{1{\bm{i}}}\cdot{\bm{S}}_{2{\bm{i}}}+\sum_{{\bm{i}}<{\bm{j}}}J_{{\bm{i}}{\bm{j}}}\,{\bm{S}}_{2{\bm{i}}}\cdot{\bm{S}}_{2{\bm{j}}}+\ldots\,.$ (2.7)
We have now explicitly included a direct exchange interaction between the ${\bm{S}}_{2{\bm{i}}}$ spins, as it will be important for our purposes below. Note that there is no Hubbard interaction $U$ in $\mathcal{H}_{\rm ALM}$, and so we have replaced the $U$ in $\mathcal{H}_{\rm ancilla-decoupled}$ by the Kondo interaction $J_{K}$ in $\mathcal{H}_{\rm ALM}$.
There are two complementary ways to connect the Hubbard model with decoupled ancillas, $\mathcal{H}_{\rm ancilla-decoupled}$, to $\mathcal{H}_{\rm ALM}$, as indicated in Fig. 7.
  • Starting from the Hubbard model, we perform a Hubbard-Stratonovich transformation on the Hubbard interaction by a vector ‘paramagnon’ field $\bm{\mathcal{P}}_{\bm{i}}$ [51]. This is an exact transformation, which replaces the Hubbard interaction in Eq. (2.6) by
@@
    This is a paramagnon theory with 3 local harmonic oscillators on each site. The ground state is when all three oscillators are in the $n=0$ state: $\left|0,0,0\right\rangle$. There is a triplet of first excited states:
@@
    We can map this low energy spectrum to that of a pair of $S=1/2$ spins with a mutual interaction $J_{\perp}{\bm{S}}_{1{\bm{i}}}\cdot{\bm{S}}_{2{\bm{i}}}$. By comparing the above matrix elements to those that couple the singlet and triplet spin states states, we obtain the operator identification
@@
    Inserting Eq. (2.10) into Eq. (2.8), we obtain $\mathcal{H}_{\rm ALM}$, but with an additional ferromagnetic Kondo interaction between the electrons $c_{{\bm{i}}\alpha}$ and spins ${\bm{S}}_{2{\bm{i}}}$. Ferromagnetic Kondo couplings are expected to be irrelevant, and we will note this coupling again in Eq. (5.3). So we see that the ancilla spins are simply the fractionalization of the familiar $S=1$ paramagnon into a pair of $S=1/2$ spins [51].
  • Starting from $\mathcal{H}_{\rm ALM}$, we perform a Schrieffer-Wolff transformation to map to $\mathcal{H}_{\rm ancilla-decoupled}$ [152]. This Schrieffer-Wolff transformation eliminates all the excited states of the ancilla spins in powers of $1/J_{\perp}$, and returns to the single band Hubbard model for the $c_{{\bm{i}}\alpha}$, with the value of $U$ shown in Eq. (2.14) and Fig. 7 [152]. At leading order in $1/J_{\perp}$, we can demonstrate this mapping by considering only a single value of $\bm{i}$}. Then at each $\bm{i}$}, the ancilla Hamiltonian is
    $H_{1}=J_{\perp}{\bm{S}}_{1}\cdot{\bm{S}}_{2}-\mu c_{\alpha}^{\dagger}c_{\alpha}+J_{K}{\bm{S}}_{1}\cdot c_{\alpha}^{\dagger}\frac{{\bm{\sigma}}_{\alpha\beta}}{2}c_{\beta}$ (2.11)
    $E_{0}$ $=$ $-\frac{3J_{\perp}}{4},~~E_{2}=-\frac{3J_{\perp}}{4}-2\mu\,$,
    $E_{1}$ $=$ $-\frac{3J_{\perp}}{4}-\mu-\frac{3J_{K}^{2}}{16J_{\perp}}+\mathcal{O}(1/J_{\perp}^{2})\,.$ (2.12)
    These states map onto those of a Hubbard model
    $H_{2}=-\mu^{\prime}c_{\alpha}^{\dagger}c_{\alpha}+U(c_{\uparrow}^{\dagger}c_{\uparrow})(c_{\downarrow}^{\dagger}c_{\downarrow})$ (2.13)
    with parameters
    $U=\frac{3J_{K}^{2}}{8J_{\perp}},\quad\mu^{\prime}=\mu+\frac{3J_{K}^{2}}{16J_{\perp}}\,.$ (2.14)
While the above derivations are valid for large $J_{\perp}$, for general $J_{\perp}$ the mapping from the ALM to the Hubbard model requires multiple gauge fields, as is reviewed in the Appendix and elsewhere [4]. The additional gauge fields are higgsed in the underdoped region, and so we need not refer to them for now; we will mention them in Section 8 below Eq. (8.6).
@@
Figure 8: Phases of the Ancilla Layer Model. The noted areas are per spin. The phases are distinguished by the condensation of the boson $\Phi$, which hybridizes the conduction electrons in the top layer with the fermionic spinons in the middle layer.
@@
Figure 9: Comparison of the phases of the Kondo lattice model in Fig. 5, and the Ancilla Layer Model of the single band Hubbard model in Fig. 8. There is an inversion in the phase in which $\Phi$ is condensed.
@@
Figure 10: Comparison between (a) the quantum dimer theory of the square lattice FL* state [126] and the (b) ancilla layer theory. The dashed green dimer is the state $c_{{\bm{i}}\alpha}^{\vphantom{\dagger}}f_{1{\bm{i}}\alpha}^{\dagger}$ created by the hybridization $\Phi$ in the ALM. Note that the ALM has exactly NN blue dimers more than the dimer model, and these extra dimers are the number in a trivial rung-singlet bilayer antiferromagnet.
@@
The connection in Fig. 10 shows that the ALM is, in a sense, ‘minimal’. For a FL* mean field theory with both a small Fermi surface and spinons, we need the middle and bottom ancilla layers to provide the vacancy-spin and singlet valence bonds respectively. And a bonus, not available in the quantum dimer analysis of Ref. [126] or other approaches [138, 144, 145], is that it also provides a mean-field theory which captures the FL large Fermi when the bottom two layers are in a rung-singlet state.
@@
The wavefunction Eq. (2.15) has been studied numerically in Refs. [53, 54, 55], with the couplings in $\mathcal{H}_{\rm KLmf}$ treated as variational parameters. The results of Refs. [53, 54] successfully capture the evolution of local, multi-point spin and charge correlations with doping as measured in cold atom experiments on the square lattice fermionic Hubbard model [56]. The results of Ref. [55] compare well with the exact diagonalization of Hubbard models in one and two dimensions.

### v6

@@ 2 Fractionalized Fermi liquids
An important step forward was Oshikawa’s proof of the Luttinger constraint using a ’t Hooft anomaly matching argument [156]. Oshikawa identified a mixed anomaly between translations and global U(1) symmetry of charge conservation. The value of the anomaly can be computed exactly from a knowledge of the density of electrons in the lattice Hamiltonian. Matching this anomaly with that of the low energy Fermi liquid theory, Oshikawa established the Luttinger constraint in a non-perturbative manner; see Ref. [7] for a review.
The idea of a ‘fractionalized Fermi liquid’ (FL*) was introduced in Refs. [12, 13], as a metallic state in which the Fermi surface did not obey the Luttinger constraint. In the simplest case, the volume enclosed by the Fermi surface in FL* was the same as free electrons with density $\rho-1$ (we are assuming a spin $S=1/2$ degeneracy). The central point [13, 56, 157, 158, 2, 3, 58, 159, 7] was that it was possible to satisfy Oshikawa’s anomaly by combining the anomaly of a Fermi surface (which contributes an amount equivalent to a density $\rho-1$) with that of a fractionalized spin liquid of the type we will study in Section 3 (which is quantized to an amount equivalent to a density 1). It is trivially possible to shift $\rho$ by an even integer by adding and removing filled bands, and the novelty is the shift in FL* by an odd integer.
@@
These ideas apply rather generally to electronic lattice models, and are summarized schematically in Fig. 3. An explicit construction of FL* is simplest and well established for the Kondo lattice model, and this will be described in Section 2.1. Our main interest here is the FL* phase in a single band Hubbard-like model, and we will turn to this in Section 2.2.
@@ 2.1 Kondo lattice
The $\Phi$ is the decoupling field for the $J_{K}$ interaction, while the $t_{1,{\bm{i}}{\bm{j}}}$ are the decoupling fields for the $J_{1,{\bm{i}}{\bm{j}}}$ interactions. Unlike Section 3.2, we do not assume any flux in the $t_{1,{\bm{i}}{\bm{j}}}$, and make the simplest possible choice in which the $t_{1,{\bm{i}}{\bm{j}}}$ have the same symmetry as the lattice; consequently, the $f_{1}$ spinons initially form a spinon Fermi surface. With these choices, our Kondo lattice theory only has an emergent U(1) gauge symmetry after adding an emergent gauge field $a_{{\bm{i}}{\bm{j}}}$ via the Peierls factor $t_{1,{\bm{i}}{\bm{j}}}\rightarrow t_{1,{\bm{i}}{\bm{j}}}e^{ia_{{\bm{i}}{\bm{j}}}}$. The gauge transformations are
@@
  • $\langle\Phi\rangle=0$, FL*. This is the novel phase, in which the $f_{1}$ and $c$ fermions are decoupled to leading order in $\mathcal{H}_{\rm KLmf}$. The $f_{1\alpha}$ form a spin liquid, in this case one with a spinon Fermi surface. But any other spin liquid with the same anomaly is allowed i.e. any other spin liquid with unit density of spinons in the ground state. The conduction electrons now form a ‘small’ Fermi surface of area $(\rho-1)/2=p/2$ (mod 1) per spin (where $\rho$ is the total density of electrons of both spins in the Kondo lattice), and this is not the Luttinger value. At higher order, the $f_{1}$ and $c$ fermions will couple with each other, but by extending the general arguments of Luttinger, we can conclude that such effects will not charge the Fermi surface area. Note that a solution with $\langle\Phi\rangle=0$ is not permitted in the single-spin Kondo model because $J_{K}$ flows to infinity under renormalization. In the lattice model, the flow of $J_{K}$ is cutoff at a scale determined by $J_{1{\bm{i}}{\bm{j}}}$ [54, 160].
@@
  • FL* in the Anderson lattice. It also interesting to consider the formation of FL* in an Anderson lattice model, where the ${\bm{S}}_{1}$ spins are replaced by electronic orbitals with a large local repulsion $U$, and the total density of electrons in both bands remains at $1+p$. In general, the Anderson lattice model is not particle-hole symmetric, and so the density of electrons in the $f_{1}$ band will deviate from unity, and consequently the actual density of electrons in the $c$ band will deviate from $p$. Nevertheless, a crucial point is that the area of the Fermi surface per spin will remain pinned at $p/2$ (mod 1) because the anomaly of the spin liquid is quantized at unity.
The existence of a FL* state in a two-band model can be extended into an essentially rigorous argument [13, 7]. We start from the decoupled limit, where the $f_{1}$ band forms a stable quantum spin liquid, and the conduction electrons form their own Fermi surface. This decoupled state is stable to the $J_{K}$ coupling because such a coupling is free of infrared divergencies and cannot destabilize the quantized anomaly of the spin liquid. Moreover, by the usual Luttinger-type analysis, the volume enclosed by the Fermi surface of the conduction electrons cannot charge to all orders in $J_{K}$. Related arguments appear in analyses based upon solvable Kitaev spin liquids [60, 61, 62].
A number of numerical studies [161, 162, 163, 164, 165, 166, 167] support the existence of phases similar to FL and FL* in various Kondo lattice models.
@@ 2.2 Single band model
We now turn to a construction of the FL* phase in a single band model, such as the square lattice Hubbard model (possibly with additional short-range interactions) of interest for the cuprates. Now the situation is more complicated than in the Kondo lattice model, as there is no natural distinction between electrons that form a spin liquid, and electrons that form a Fermi surface. Nevertheless, strong physical arguments were made for the existence of FL* even in a single band model [69, 70, 2, 71, 72, 73, 74, 75]. Numerical studies of the Hubbard model have not provided complete evidence for the FL* phase, but given the small size of the pockets in the Yamaji experiment, and the limited resolution of all numerical studies, this is perhaps not surprising. We will briefly review the theoretical arguments, and then turn to the ALM as the best current method of obtaining the FL* state in a single band model.
At the level of cartoon pictures, we can illustrate the structure of a FL* state in Fig. 6 [74]. This figure describes three distinct metallic phases in a Hubbard time model with electron density 1p1-p. Recall that in the absence of a broken symmetry, the Luttinger constraint on hole Fermi surfaces is that they have a fractional area per spin of $(1+p)/2$, relative to the area of the full square lattice Brillouin zone.
Figure 6: Cartoon pictures of different states of doped antiferromagnets; adapted from Ref. [74]. The areas are those that would be measured by a probe such as quantum oscillation. The AF metal has long-range antiferromagnetic order, and the reduced Brillouin zone is shown with the dashed line. The other phases do not break any symmetries. The open circles are holons, and these are assumed fermionic in the holon metal. The green dimers represent bound states of holons and spinons. The holon metal pockets do not have any symmetry associated with the reduced Brillouin zone of the AF metal. All 3 states have a Hall co-efficient of density $p$ positively charged carries.
@@
  • Holon metal. This is a state with no broken symmetry, in which the electrons have paired up in singlet bonds which resonate with each other. The dopants are realized by spinless mobile vacancies of charge +e+e, known as holons. The density of holons is $p$, and if the holons are fermions, they will form Fermi surfaces corresponding to spinless free fermions of density $p$. If there are four distinct Fermi surfaces in the Brillouin zone (as is the case in many computations), then the fractional area of each pocket will be $p/4$. Although this area is the same as that for the AF metal, the reason is very different. Now there is no broken symmetry, and the fermionic quasiparticles are spinless holons. This Fermi surface area does not obey the Luttinger constraint, and this is permitted because of the presence of the spin liquid. The holon metal pockets do not have reflection symmetry about the reduced Brillouin zone boundary shown for the AF metal. A systematic, gauge theoretic treatment of the holon metal state is available in the literature [168, 169, 170, 171, 172, 173, 174, 175, 176, 177], including a previous review by one of us [178], and applications to the Lieb lattice Hubbard model are discussed in Refs. [179, 180]. We note in passing that with the generalized definition in Fig. 3, the holon metal is also a type of FL*.
  • FL*. Finally, we turn to the metallic state of interest. The holon metal state also has spinon excitations, which can be created out of the ground state. Now imagine a situation in which each holons gains energy by binding with a spinon, so that the system can pay the price for creating the spinons. Then the ground state will change into one in which the mobile charge carriers are holon-spinon bound states [181, 69, 70, 2, 182, 71, 183, 72, 73, 74, 75, 184, 185, 186, 187, 188]. These bound states are always fermions with charge +e+e, spin $S=1/2$, just like a hole. Treating these holes as free fermions, we conclude that the total area of the Fermi surface should be $p/2$. If there are 4 distinct pockets (as there in the computation below), then each pocket will have the distinctive area of $p/8$. This Fermi surface area also does not obey the Luttinger constraint.
As we noted in Section 1.1, and will discuss further in Section 5, recent observations by Chan et al. [10] of the Yamaji effect in the cuprate HgBa2CuO4+δ show a fractional area of ‘approximately $1.3\%$’ at a doping $p=0.1$, close to the value $p/8=1.25\%$ predicted for FL* [12, 13, 69, 2, 189, 78]. Moreover, the Yamaji effect requires coherent interlayer transport. This is not possible for the holon metals as the holons carry charges of distinct emergent gauge fields within each layer, but the FL* quasiparticles are gauge neutral and can indeed tunnel coherently between layers. In the AF metal coherent tunneling requires significant interlayer spin correlations, which are not observed [80, 81]. All of these facts provide strong evidence in favor a FL* description of the pseudogap phase of the cuprates [78].
@@ 2.2.1 Layer construction with ancilla qubits
We wish to have a mean-field theory which changes a large hole-like Fermi surface of area $(1+p)/2$ to small hole-like Fermi surfaces of total area $p/2$. The simplest way to achieve this in mean-field theory is to hybridize the large Fermi surface with another band at half-filling (as in the Kondo lattice, and was done by Yang-Rice-Zhang (YRZ) [190, 79])—this leads to a Fermi surface of the needed area $(1+p+1)/2$ (mod 1) =p/2=p/2. But we are not allowed to add a single band at half-filling because it is not ‘trivial’ i.e. it is not anomaly-free, and its excitations cannot be integrated out because the extra band can only acquire gap with a broken symmetry. On the other hand, we are allowed to add two bands at half filling each (or any even number of bands), because they can form a trivial insulator with a gap. After the first added band hybridizes with the physical electrons to yield the small Fermi surfaces, the second added band is left decoupled, and it must form a spin liquid to satisfy the Oshikawa anomaly. This is the essence of the ALM.
We begin with a constructive derivation of the ALM starting from the familiar Hubbard model
$\mathcal{H}_{\rm Hubbard}$ $=$ $-\sum_{{\bm{i}},{\bm{j}}}t_{{\bm{i}}{\bm{j}}}c^{\dagger}_{{\bm{i}}\alpha}c_{{\bm{j}}\alpha}+U\sum_{\bm{i}}(c_{{\bm{i}}\uparrow}^{\dagger}c_{{\bm{i}}\uparrow})(c_{{\bm{i}}\downarrow}^{\dagger}c_{{\bm{i}}\downarrow})\,.$ (2.6)
Initially, we employ the familiar mapping of the Hubbard model to a theory of paramagnons coupled to the Fermi surface. We perform a Hubbard-Stratonovich transformation on the Hubbard interaction by a vector ‘paramagnon’ field $\bm{\mathcal{P}}_{\bm{i}}$ [77]. This is an exact transformation, which replaces the Hubbard interaction in Eq. (2.6) by
@@
This is a paramagnon theory with 3 local harmonic oscillators on each site. Now we take steps different from conventional paramagnon approaches. The ground state of the paramagnon theory is obtained when all three oscillators are in the $n=0$ state: $\left|0,0,0\right\rangle$. There is a triplet of first excited states:
@@
We can map this low energy spectrum to that of a pair of $S=1/2$ ancilla spins with a mutual interaction $J_{\perp}{\bm{S}}_{1{\bm{i}}}\cdot{\bm{S}}_{2{\bm{i}}}$. By comparing the above matrix elements to those that couple the singlet and triplet spin states states, we obtain the operator identification
@@
Inserting Eq. (2.9) into Eq. (2.7), we obtain the Hamiltonian of the ALM, $\mathcal{H}_{\rm ALM}$, which is a simple augmentation of the Kondo lattice Hamiltonian $\mathcal{H}_{\rm KL}$ in Eq. (2.1) by the layer of ${\bm{S}}_{2{\bm{i}}}$ spins
$\mathcal{H}_{\rm ALM}=\mathcal{H}_{\rm KL}+J_{\perp}\sum_{{\bm{i}}}{\bm{S}}_{1{\bm{i}}}\cdot{\bm{S}}_{2{\bm{i}}}+\sum_{{\bm{i}}<{\bm{j}}}J_{{\bm{i}}{\bm{j}}}\,{\bm{S}}_{2{\bm{i}}}\cdot{\bm{S}}_{2{\bm{j}}}\,.$ (2.10)
This Hamiltonian is illustrated in the bottom half of Fig. 7.
Figure 7: Illustration of the mapping from a single-band Hubbard model with decoupled ancilla qubits, to a single band with free fermions coupled to a bilayer antiferromagnet. The Schrieffer-Wolff transformation is derived in Ref. [191] and reviewed in Appendix A, and the Hubbard-Stratonovich transformation is derived in Ref. [77]. The layer numbers of the layer construction with ancilla qubits are indicated in the bottom picture; these appear in Tables 2 and 3.
So we see that the ancilla spins are simply the fractionalization of the familiar $S=1$ paramagnon into a pair of $S=1/2$ spins [77]. The above derivation of the mapping to the ALM yields an additional ferromagnetic Kondo interaction between the electrons $c_{{\bm{i}}\alpha}$ and spins ${\bm{S}}_{2{\bm{i}}}$. Ferromagnetic Kondo couplings are expected to be irrelevant, and we will note this coupling again in Eq. (5.3). We have also explicitly included a direct exchange interaction between the ${\bm{S}}_{2{\bm{i}}}$ spins, as it will be important for our purposes below. Note that there is no Hubbard interaction $U$ in $\mathcal{H}_{\rm ALM}$, and so we have replaced the $U$ in $\mathcal{H}_{\rm ancilla-decoupled}$ by the Kondo interaction $J_{K}$ in $\mathcal{H}_{\rm ALM}$. A complementary mapping from the ALM to the Hubbard model via the Schrieffer-Wolff transformation, also indicated in Fig. 7, is described in Appendix A.
While the derivations of $\mathcal{H}_{\rm ALM}$ are valid for large $J_{\perp}$, for general $J_{\perp}$ the mapping from the ALM to the Hubbard model requires multiple gauge fields, as is reviewed in Appendix B and elsewhere [7]. The additional gauge fields are higgsed in the underdoped region, and so we need not refer to them for now; we will mention them in Section 8 below Eq. (8.6).
@@
Figure 8: Phases of the Ancilla Layer Model. The noted areas are per spin. The phases are distinguished by the condensation of the boson $\Phi$, which hybridizes the conduction electrons in the top layer with the fermionic spinons in the middle layer. The ${\bm{S}}_{2}$ spin liquid will be described in Section 3.
@@
We can now summarize the FL* construction of the Hubbard model in the following simple terms. We express the Hubbard model as a theory of antiferromagnetic paramagnons, and then fractionalize the $S=1$ paramagnon into two $S=1/2$ spins, ${\bm{S}}_{1}$ and ${\bm{S}}_{2}$. The ${\bm{S}}_{1}$ spins form a well-established Kondo lattice heavy Fermi liquid with the electrons $c_{\alpha}$ to yield the hole pockets. (We emphasize that it is the FL phase of the Kondo lattice which maps to the FL* phase of the ALM, as also noted in Fig. 9.) The ${\bm{S}}_{2}$ spins form the spin liquid needed to satisfy the Oshikawa anomaly. Note that the ${\bm{S}}_{2}$ spins can form any spin liquid appropriate for $S=1/2$ spins, and we will not choose the spinon Fermi surface state for the bottom layer; instead, the suitable spin liquid will be introduced in Sections 3.
@@
Figure 9: Comparison of the phases of the Kondo lattice model in Fig. 5, and the ALM of the single band Hubbard model in Fig. 8. There is an inversion in the phase in which $\Phi$ is condensed: the well-established, $\Phi$-condensed, ‘large’ Fermi surface phase of the Kondo lattice maps to the FL* pseudogap phase of the ALM.
@@
Figure 10: Schematic comparison between (a) the quantum dimer theory of the square lattice FL* state [74] and the (b) ALM. The dashed green dimer is the state $c_{{\bm{i}}\alpha}^{\vphantom{\dagger}}f_{1{\bm{i}}\alpha}^{\dagger}$ created by the hybridization $\Phi$ in the ALM. Note that the ALM has exactly NN blue dimers more than the dimer model, and these extra dimers are the number in a trivial rung-singlet bilayer antiferromagnet.
@@
The connection in Fig. 10 shows that the ALM is, in a sense, ‘minimal’. For a FL* mean field theory with both a small Fermi surface and spinons, we need the middle and bottom ancilla layers to provide the vacancy-spin and singlet valence bonds respectively. And a bonus, not available in the quantum dimer analysis of Ref. [74, 75] or other approaches [181, 72, 73], is that it also provides a mean-field theory which captures the FL large Fermi when the bottom two layers are in a rung-singlet state.
@@
The wavefunction Eq. (2.11) has been studied numerically in Refs. [82, 83, 84], with the couplings in $\mathcal{H}_{\rm KLmf}$ treated as variational parameters. The results of Refs. [82, 83] successfully capture the evolution of local, multi-point spin and charge correlations with doping as measured in cold atom experiments on the square lattice fermionic Hubbard model [85, 86, 87]. The results of Ref. [84] compare well with the exact diagonalization of Hubbard models in one and two dimensions.